A sample of N2O contains 3.75×1023 molecules. How many moles of N2O are present?

Find the mass of the sample in Part 1. 


1
Expert's answer
2021-04-05T07:42:02-0400

no. of moles of N2O present = "\frac{3.75×10^{23}}{6.022×10^{23}}"

= 0.622 mol


Mass of N2O = 0.622 × 44.013 = 27.376g